40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 338     Entry time: Fri Apr 7 16:19:03 2017
Author: Gabriele 
Type: Optics 
Category: Daily Progress 
Subject: New 1103P laser installed 

This afternoon I installed the new Lumentum (former JDSU) HeNe laser, model 1103P in CR0.

Realigned everything.

I installed a sample in the chamber to reflect a beam back inot the QPD. Checking the QPD signals over a hour and more did not show any sign of excess noise or instability.

ELOG V3.1.3-